The diameter of a brand of Ping Pong balls is approximately normally distributed, with a mean of 1.31 inches and a standard deviation of 0.08 inches. A random sample of 4 balls is selected.
-The probability is 50% that the sample mean will be between what two values, symmetrically distributed around the population mean?
Lower bound: __ inches
Upper bound: __ inches
I'm pretty sure I need to start by finding a Z score, but am having a hard time figuring which formulas I need to use, for that and to finish the problem. Thanks.

Answer by stanbon(75887)

The diameter of a brand of Ping Pong balls is approximately normally distributed, with a mean of 1.31 inches and a standard deviation of 0.08 inches. A random sample of 4 balls is selected.
-The probability is 50% that the sample mean will be between what two values, symmetrically distributed around the population mean?
-----
Looks like you want a 50% confidence interval.
----
sample mean = x-bar = 1.31
Margin of Error = ME = z*s/sqrt(n))= 0.6745*0.08/sqrt(4) = 0.0270
------
50%CI:: 1.31-0.0270 < u < 1.31+0.0270
---------------
Cheers,
Stan H.
--------------
Lower bound: 1.2831 inches
Upper bound: 1.3369 inches